OpenStudy (anonymous):

What is the 8th term of the geometric sequence where a1 = 256 and a3 = 16?

OpenStudy (amistre64):

divide off the 256 to get a1 to a 1 and the rest should fall into suit

OpenStudy (amistre64):

in other words: \[r = \sqrt{\frac{16}{256}}\] \[a8 = r^7\]

OpenStudy (anonymous):

i still dont get it srry

OpenStudy (amistre64):

there are afew underlying concepts to understand; youd have to dialogue with me in order for me to see where it is you need the help

OpenStudy (dumbcow):

use the formula \[a_{n} = a_{1}*r^{n-1}\]

OpenStudy (amistre64):

o yeah, i forgot about the 256 i lopped off :)